OpenStudy (anonymous):

if MN is the diameter of the circle, with M (-3,-4) and N (-7,-14) find the center of the circle

OpenStudy (johnweldon1993):

Well you know that halfway of the diameter is the center...so we just need to find the mid point... For the x-coordinate \[\large \frac{ x_1 + x_2 }{ 2 }\] \[\large \frac{ -3 - 7 }{ 2 }\] \[\large \frac{ -10 }{ 2 }\] \[\large -5 \] and for the y-coordinate \[\large \frac{ y_1 + y_2 }{ 2 }\] \[\large \frac{ -14 - 4 }{ 2 }\] \[\large \frac{ -18 }{ 2 }\] \[\large -9\] so your center is..\(\large (-5 , -9) \)